A sample of potassium phthalate, KHC8H4O4, monobasic, weighing 4.0700g is titrated with NaOH solution and back titrated with HCl. NaOH required = 46.40 mL; HCl required = 5.35mL. If each mL of the HCl is equivalent to 0.01600g of Na2O, what volume of H2O or of 6.00N NaOH must be added to 500mL of the above NaOH to bring it to 0.5000N?
1
Expert's answer
2020-11-26T05:22:52-0500

2HCl(aq)+Na2O(s)=H2O(l)+2NaCl(aq)

Moles of Na2O= 0.016/62=.000258moles

Moles of HCl= 0.000258×2

                     =0.000516moles

Total moles of acid in 5.35 ml= 5.35×0.000516

                                                =0.0027606moles

HCl(aq)+NaOH(aq)=H2O(l)+NaCl(aq)

Moles of NaOH= 0.0027606moles

Molarity=1000×0.00027606/46.4

            =0.0595M

C1V1=C2V2

46.40×0.0595=500C2

C2= 0.0055216M

Molarity= Normality/Equivalent

For 0.5000N NaOH=0.5/1

                               =0.5M

For 6.000M NaOH= 6/1 =6M

C1V1=C2V2

C1= (6+0.0055216) =6.0055216M

V1= 500mL

C2= 0.5M

V2=?

6.0055216×500=0.5 V2

V2= 6005.5216mL

   =6.006L

Volume of water added= (6.006-0.5) = 5.506L